OpenStudy (anonymous):

Solve the equation:

OpenStudy (anonymous):

\[\frac{ 3x }{ x-5 }=\frac{ 2 }{ x+1}+\frac{ 2x^2+40 }{ x^2-4x-5 }\]

OpenStudy (anonymous):

I know that you have to factor or whatever the x^2-4x-5

OpenStudy (anonymous):

so it would look like this \[\frac{ 2(x^2+20) }{ (x-5)(x+4)}\]

OpenStudy (dumbcow):

x^2 -4x -5 = (x-5)(x+1)

OpenStudy (dumbcow):

\[\frac{3x}{x-5} = \frac{2(x-5) + 2x^2 + 40}{(x-5)(x+1)}\] so \[3x(x+1) = 2(x-5) +2x^2 +40\] combine terms set equal to zero solve quadratic using factoring or quadratic formula

OpenStudy (anonymous):

x=6, -5? @dumbcow

OpenStudy (dumbcow):

check your signs ... i get -6,5 but x can't equal 5 because it makes denominator 0 in original equation

OpenStudy (anonymous):

so its just -6

OpenStudy (anonymous):

@dumbcow thank you for helping me out with this one.

OpenStudy (dumbcow):

yw